OpenStudy (ronaldo7):

How many liters of pure water must be added to 300 liters of a 30% acid solution to make an acid solution that is 20%?

OpenStudy (ronaldo7):

I need help please this is due tommorow.

OpenStudy (anonymous):

does it give a formula to use

OpenStudy (anonymous):

so the amount of water in 30% acid solution is=300(1-30/100)=210 liter the amount of water in 20% acid solution is=300(1-20/100)=240 liter so the amount of water to be added is=240-210=30 liter

OpenStudy (ronaldo7):

So how do you set it up that is all I need

OpenStudy (anonymous):